# GraphToxin: Reconstructing Full Unlearned Graphs from Graph Unlearning

### v1

@@ 3.2.3 Feature Smoothness
Most real-world graphs follow the homophily principle, where neighboring nodes tend to share similar features. Graph Dirichlet energy [smooth_sparse, graph_mia, unlearning_inversion_attacks_graph] explicitly encodes and quantifies this feature smoothness:
@@ 3.2.4 Loss Function
In summary, the final objective function of GraphToxin is as follows:
@@
where the Fisher coefficient $\alpha_{1}$ and the Laplacian coefficient $\alpha_{2}$ control the fine-grained curvature matching and feature smoothness, respectively.
3.3 Extensions: Multiple Node Removal and Black-box GraphToxin
3.3.1 Multiple Node Removal
For multiple node removals, the topology of the full unlearned graph $\mathcal{G}_{d}$ is no longer a simple star form. Instead, its topology $\mathcal{E}_{\mathcal{N}_{\mathcal{V}_{d}^{1}}}$ becomes significantly more intricate due to the interconnection of removed nodes and their affected neighbors.
To simultaneously reconstruct node features and the more complex graph topology, we first initialize the graph topology by sampling a fixed number of edges using Bernoulli distribution [gradient_attack_gnn]. This initialization is defined as $A_{0}\sim Bernoulli(p)$, where $p=\frac{|\mathcal{E}_{\mathcal{N}_{\mathcal{V}_{d}^{1}}}|}{\binom{|\mathcal{V}_{{rec}_{\mathcal{V}_{d}}}|}{2}}$ controls the expected edge density and statistically aligns with the known edge budget. Next, we adopt the same final objective function $\mathfrak{L}_{white}$ used for a single node removal, but we update the adjacency matrix $A_{\theta}$ with the projected gradient method. Following Sinha et al. [gradient_attack_gnn], each entry $(i,j)$ of $\tilde{A}$ is updated through the entry-wise projection operator as:
@@ 3.3.1 Multiple Node Removal
Finally, we repeat the sampling process using the learned edge probabilities from the last iteration to generate the final reconstructed adjacency matrix $\hat{A}$.
@@ 3.3.2 Black-box GraphToxin
Distinct from the white-box setting, the black-box GraphToxin cannot access gradients and only relies on posterior probabilities. Notably, this most constrained attack remains feasible in reality. Inspired by Zhuang et al. [data_free_msa_gnn], attackers first steal the original GNN $\mathcal{F}_{ori}$ and unlearned GNN $\mathcal{F}_{un}$ using a data-free and black-box model extraction attack, which involves a graph generator $\phi_{G}(\cdot)$ and a surrogate model $\phi_{S}(\cdot)$. Specifically, attackers employ the zeroth-order gradient estimation [data_free_msa, data_free_msa_gnn] to approximate the gradients of $\mathcal{F}_{ori}$ and $\mathcal{F}_{un}$:
$\tilde{\nabla}\mathcal{L}(\theta_{\phi_{G}{(\cdot)}},\ddot{\mathcal{G}},u_{i})=\frac{d(\mathcal{L}(\theta_{\phi_{G}{(\cdot)}},\ddot{\mathcal{G}}+u_{i})-\mathcal{L}(\theta_{\phi_{G}{(\cdot)}},\ddot{\mathcal{G}})}{\epsilon}u_{i}$ (12)
where the generated graph $\ddot{\mathcal{G}}=\phi_{G}(\theta_{\phi_{G}},z)$; the Gaussian noise $u_{i}\sim\mathrm{N}(0,\mathrm{I}_{\mathcal{D}})$, and $\cdot$ can be $\mathcal{F}_{ori}$ or $\mathcal{F}_{un}$, depending on the stealing target. $\mathcal{D}$ is the feature dimension and $\epsilon$ is a small positive constant called the smoothing factor. However, Equation (12) results in biased estimation as it only involves a single random direction. Following previous studies [data_free_msa, data_free_msa_gnn], the average gradient estimation from $m$ different random directions is required for precise gradient estimation, which is denoted as:
$\tilde{\nabla}\mathcal{L}(\theta_{\phi_{G}{(\cdot)}},\ddot{\mathcal{G}})=\frac{1}{m}\sum\limits_{i=1}^{m}\tilde{\nabla}\mathcal{L}(\theta_{\phi_{G}{(\cdot)}},\ddot{\mathcal{G}},u_{i})$ (13)
We adopt the same architecture and full parameterization trick as Zhuang et al. [data_free_msa_gnn] to generate the query graph $\ddot{\mathcal{G}}$. To encourage the generator $\phi_{G}(\cdot)$ producing more diverse and complex $\ddot{\mathcal{G}}$, $\phi_{G}(\cdot)$ is trained to maximize the disagreement between the posteriors of the victim and the surrogate models. Meanwhile, the surrogate model $\phi_{S}(\cdot)$ is trained using the query responses of $\ddot{\mathcal{G}}$ from the victim model. For details, we refer the interested readers to Zhuang et al. [data_free_msa_gnn].
Once attackers obtain the copycats $\tilde{\mathcal{F}}_{ori}$ and $\tilde{\mathcal{F}}_{un}$, they exploit the gradient difference between these copycats with the same $\mathfrak{L}_{white}$ for full unlearned graph recovery. However, the reconstruction quality highly depends on the stealing quality, and the gradient difference induced by two copycats inevitably introduces noises. To mitigate these issues, we add the following loss as the semantic calibration in $\mathfrak{L}_{white}$:
$\mathfrak{L}_{seman}=CE(\tilde{\mathcal{G}_{d}},Y_{d})=-\sum_{i=1}^{|\mathcal{V}_{{rec}_{\mathcal{V}_{d}}}|}Y_{d_{i}}\cdot\log p_{d_{i}}$ (14)
where $CE$ is the cross-entropy loss to minimize the discrepancies between the ground-truth labels and the queried predictions of the recovered nodes in the synthetic graph $\tilde{\mathcal{G}_{d}}$.
Therefore, the final objective $\mathfrak{L}_{black}$ of the black-box GraphToxin is updated as:
@@
where $\alpha_{3}$ is the calibration coefficient.
@@ 3.4 Evaluation Framework Design
To answer the third key research question, we propose a systematic and comprehensive evaluation framework for GraphToxin. We will further demonstrate its necessity along with the worst-case evaluation in Experimental Validation.
@@ 3.4.1 Feature-level Evaluation
Root Normalized Mean Squared Error (RNMSE) Following Sinha et al. [gradient_attack_gnn], we leverage RNMSE to measure the discrepancies between the recovered node features and their corresponding ground-truth node features.
$\text{RNMSE}(X_{v},\tilde{X}_{v})=\frac{\|X_{v}-\tilde{X}_{v}\|}{\|X_{v}\|}$ (16)
For each $v\in\mathcal{V}_{{rec}_{\mathcal{V}_{d}}}$, we compute its RNMSE to the corresponding node features in the original training graph $\mathcal{G}$ and then report the average RNMSE of all nodes in $\mathcal{V}_{{rec}_{\mathcal{V}_{d}}}$.
@@ 3.4.2 Global-level Evaluation
For a single node removal, the graph topology is fixed in GraphToxin. There is no need to introduce structural-level evaluation. While for multiple node removals, traditional metrics like Area Under Curve (AUC) or Average Precision (AP) focus only on the existence of individual links and fail to capture the holistic structural patterns. Instead, we provide the following global-level metrics to jointly evaluate the full feature and topology recovery.
Embedding Distance (ED) During message passing, node embedding inherently encodes node features, and local and high-level structural patterns. Consequently, quantifying the similarity between the node embeddings of the recovered and ground-truth deleted graphs provides a more holistic view compared to single-modality metrics. We consider the squared quadratic Wasserstein distance (p=2) [vaserstein1969markov] to measure this similarity, denoted as $W_{2}(\tilde{H}^{k},H^{k})$.
Pyramid Match Graph Kernel (PMGK) PMGK is a robust measure for graph matching [pmgk]. It can capture structural similarities across multiple scales, from local motifs to global community structures. Specifically, we discretize the continuous embeddings of the recovered and ground-truth graphs into a set of discrete labels via K-means clustering, thus grouping nodes with similar structural or semantic patterns. Finally, we create a multi-resolution histogram of these labels of each graph and then take a weighted sum of the intersections of these two histograms. In our experiment, we utilize the grakel library [grakel] to compute PMGK.
@@ 3.4.3 Performance-level Evaluation
Intuitively, if two graphs are similar and fed into the same model, their outputs should also be similar. We adopt the following performance-level metrics to assess this alignment.
Attack Accuracy (ATT. ACC) Attack accuracy quantifies how effectively GraphToxin recovers the functionally equivalent full unlearned graph from graph unlearning.
Attack Fidelity (ATT. FID) Attack fidelity quantifies how accurately the recovered graph can yield the same model-specific predictions as the full ground-truth unlearned graph.
Posterior Wasserstein Distance (PWD) PWD measures how closely the recovered graph replicates the confidence and uncertainty of the ground-truth unlearned graph. Similarly, we adopt the squared quadratic Wasserstein distance (p=2) to compare these two posterior probabilities.
@@ 4 Experimental Validation
In this section, we conduct empirical experiments on diverse GNN variants, graph unlearning methods and graph datasets to answer our three primary research questions. Specifically, we investigate the following key subquestions:
@@
SQ2: How GraphToxin performs for a single node removal?
@@
SQ6: How GraphToxin performs in multiple node removals?
@@ 4.1.1 Datasets
To assess the effectiveness of GraphToxin, we conduct experiments on three publicly available graph datasets: Cora, PubMed [cora_pub_data], and Amazon-Photo (Photo) [amazon_photo]. We provide the data details in Appendix .1.1.
@@ 4.1.2 Unlearning Methods
We consider the following graph unlearning methods:
  1. 1.
    Exact Unlearning: After node removals, we directly retrain the GNNs from scratch (Retraining).
  2. 2.
    Approximate Unlearning: We consider two state-of-the-art graph unlearning methods with distinct frameworks: MEGU [MEGU] and ETR [ETR]. MEGU is a learning-based method that employs predictive and unlearning modules to mitigate unlearning effects, while ETR is a training-free method designed to erase the influence of deleted nodes and rectify the collateral damage to their neighbors.
@@ 4.1.4 Baselines
Since GraphToxin is the first graph reconstruction attack against graph unlearning, there is no reliable existing method to compare. Instead, we consider:
  1. 1.
    Random Graph (Rand.): We generate random graphs using Gaussian noises without parameter updates.
  2. 2.
    Learning with A Few Epochs (FewE.): We train GraphToxin for only the top $1\%$ epochs.
@@ 4.1.5 Implementation
All experiments are implemented on $4$ Nvidia A100100 GPUs. For detailed hyperparameter settings and architecture choices, please refer to Appendix .1.2.
@@ 4.2 Attack Performance (SQ1 - SQ3)
The original training and graph unlearning results with different GNN backbones across three graph datasets are detailed in Table V in Appendix .1.3, which confirm the effectiveness and consistency of all considered graph unlearning methods. Specifically, all graph unlearning methods achieve very similar performance on three GNN backbones, with less than $1\%$ variation. The GCN backbone is the best attack target (the highest performance) in Cora and PubMed while SGC excels in Photo, with less than $2\%$ variation from GCN’s results. Among the unlearning methods, ETR achieves superior performance in Cora and PubMed, while Retraining performs better with GCN and SGC backbones, and MEGU exceeds others with GraphSAGE in Photo. For the removal settings, we observe that removing $10\%$ of the highest-degree nodes (the worst-case scenario) results in larger performance degradation compared to the random node removal. This degradation is amplified for sparse and smaller-scale graphs, highlighting the necessity of a rigorous worst-case evaluation for graph unlearning.
4.2.1 Necessity of Comprehensive Evaluation
@@
We first motivate the necessity for a comprehensive evaluation framework before presenting our detailed attack analysis. We contend that the standard NRMSE metric in existing graph recovery attacks [gradient_attack_gnn] is insufficient on its own for assessing the reconstruction quality of node features. As shown in Table II, even a randomly generated graph can achieve a deceptively low NRMSE while failing on crucial performance-level metrics such as attack accuracy, fidelity, and posterior Wasserstein distance. This obscurity highlights how NRMSE can conceal significant qualitative differences between a recovered graph and the ground-truth, which are further amplified by the message-passing paradigm and ultimately projected as a large semantic gap. Hence, more sensitive metrics are required to discern subtle variations. Specifically, when NRMSE values are nearly identical, embedding distance can reveal a significant divergence between graphs. Moreover, since a single node removal induces a fairly small gradient change, GraphToxin achieves consistent performance across various unlearning methods. Notably, all results are different, but some variations cannot be reflected on the scale of 1e21e\!-\!2. However, PMGK can reflect such minimal changes that are imperceptible to other metrics.
4.2.2 Effectiveness
As shown in Table II, GraphToxin consistently and significantly outperforms the two baselines across all experimental settings in nearly every reconstruction quality and performance-level evaluation metrics. The superiority of GraphToxin generalizes across distinct graph datasets and successfully targets various graph unlearning methods. Meanwhile, GraphToxin’s overall advantage holds under both random and worst-case node removal scenarios, highlighting its robustness. Notably, GraphToxin achieves up to $\mathbf{11.14\times}$ better performance than the random generation method on the performance-level metrics in Cora, and up to $\mathbf{86.73\times}$ better in Photo. However, averaging performance across recovered graphs from the random node removal can be misleading, motivating the need for a worst-case analysis. PubMed exemplifies this problem. Due to its high sparsity and small label space (with only three labels), even a randomly generated graph can yield deceptively strong results on the performance-level metrics, as the random node removal often results in simpler recovery targets. This effect can artificially inflate the average result, offsetting the random baseline’s weaker performance on more complex and larger-scale graphs. In contrast, a worst-case evaluation targeting the top $10\%$ highest-degree nodes removes this masking effect, revealing a substantial decrease in the random graph’s performance and thus providing a more accurate assessment of GraphToxin’s efficacy.

### v2

@@ 3.2.3 Feature Smoothness
Most real-world graphs follow the homophily principle, where neighboring nodes tend to share similar features. Graph Dirichlet energy [smooth_sparse, graph_mia, unlearning_inversion_attacks_graph] explicitly captures and quantifies this feature smoothness:
@@ 3.2.4 Loss Function
In summary, the final objective of GraphToxin is:
@@
where $\alpha_{1}$ weights the curvature-aware loss while $\alpha_{2}$ controls feature smoothness regularization. We refer to them as the Fisher coefficient and the Laplacian coefficient, respectively.
3.3 Extensions: Multiple-Node Removal and Black-box GraphToxin
3.3.1 Multiple-Node Removal
For multiple-node removal, the topology of the full unlearned graph $\mathcal{G}_{u}$ is significantly more complex due to the interconnection of removed nodes and their affected neighbors.
To simultaneously reconstruct node features and such more complicated graph topology, we first initialize the topological structure by sampling a fixed number of edges using Bernoulli distribution [gradient_attack_gnn], marked as $\tilde{A}_{0}\sim Bernoulli(p)$, where $p=\frac{|\mathcal{E}_{u}|}{\binom{|\mathcal{V}_{u}|}{2}}$ controls the expected edge density and aligns with the known edge budget. Next, we adopt the same final objective $\mathfrak{L}_{white}$ as in the single-node removal case, while additionally updating the adjacency matrix $\tilde{A}$ via projected gradient descent [gradient_attack_gnn]. Each entry $(i,j)$ of $\tilde{A}$ is updated through the entry-wise projection operator as:
@@ 3.3.1 Multiple-Node Removal
We then repeat the sampling process using the learned edge probabilities from the last iteration to generate the final reconstructed adjacency matrix $\hat{A}$.
@@ 3.3.2 Black-box GraphToxin
Distinct from the white-box setting, black-box GraphToxin cannot access gradients and relies solely on posterior probabilities. Despite this constraint, the attack remains feasible in practice. The adversary first obtains surrogate models of the original and unlearned GNNs ($\mathcal{F}_{o}^{\prime}$ and $\mathcal{F}_{u}^{\prime}$) via an existing data-free, black-box model extraction attack [data_free_msa_gnn]. We refer the interested readers to [data_free_msa_gnn] for additional details.
Once the copycats are stolen, GraphToxin exploits the gradient difference between these copycats using the same objective $\mathfrak{L}_{white}$ for full unlearned graph recovery. However, the reconstruction quality highly depends on the stealing quality. Imperfect model extraction may introduce additional noise into the estimated gradients. We incorporate a semantic calibration loss to mitigate this effect:
$\mathfrak{L}_{seman}=CE(\tilde{Y}_{u},Y_{u})=-\sum_{i=1}^{|\mathcal{V}_{u}|}Y_{u_{i}}\cdot\log p_{u_{i}}$ (9)
where $CE$ is the cross-entropy loss to minimize the discrepancies between the actual node labels and the model predictions of the dummy graph $\tilde{\mathcal{G}}_{u}$ from the copycat $\mathcal{F}_{o}^{\prime}$.
The final objective for black-box GraphToxin is:
@@
where $\alpha_{3}$ controls the contribution of semantic calibration, referred as the calibration coefficient.
@@ 3.4 Evaluation Framework Design
3.4.1 Worst-case Scenario
Existing graph unlearning studies evaluate unlearning effectiveness under random node removals, but fail to examine worst-case scenarios where highly influential nodes are removed. These scenarios serve as a stress test rather than a favorable condition for the attack. Removing influential nodes indeed induces larger parameter changes, while simultaneously increases reconstruction difficulty. Such nodes typically possess high degrees, amplifying multi-hop dependencies and thereby complicating the attribution of entangled features and topologies. Moreover, worst-case settings align with realistic attacker incentives. Nodes with high influence, such as public figures in a social network, are more likely to trigger unlearning requests and may attract stronger attack interests than normal users.
For single-node removal, we unlearn the top 10% nodes with the highest degrees from the training graph and report their average attack results to assess the worst-case scenarios, while in multiple-node removal, the top 25, 50, 75 nodes with the highest degrees are selected and partitioned into 5 non-overlapping pairs based on their degree ranking.
@@ 3.4.2 Feature-level Evaluation
Feature-level evaluation measures the quality of node attributes within the recovered unlearned graph $\tilde{\mathcal{G}}_{u}$.
Root Normalized Mean Squared Error (RNMSE) We follow Sinha et al. [gradient_attack_gnn] and leverage RNMSE to quantify the discrepancies between the recovered and actual node features.
$\text{RNMSE}(\tilde{X}_{v},X_{v})=\frac{\|\tilde{X}_{v}-X_{v}\|}{\|X_{v}\|}$ (11)
We finally report the average RNMSE of all nodes in $\tilde{\mathcal{G}}_{u}$. While RNMSE is a standard metric for feature recovery, it does not fully capture structural similarity or functional consistency, as we empirically demonstrate in Section 4.3.1.
@@ 3.4.3 Global-level Evaluation
Global-level evaluation assesses the preservation of joint feature–topology structure of the recovered unlearned subgraph.
We assume that the adversary knows the size of the full unlearned graph. For single-node removal, the graph topology is fixed, and thus topology-level evaluation is unnecessary. For multiple-node removal, traditional link-wise reconstruction metrics like Area under Curve (AUC) and Average Precision (AP) fail to capture holistic structural patterns. Therefore, we adopt the following global-level evaluation metrics:
Embedding Distance (ED) Node embeddings encode both node features and multi-hop structural patterns. Consequently, we measure the similarity between the embeddings of the recovered and ground-truth unlearned graphs, which are both extracted from the original GNN model, denoted as $\tilde{H}^{k},H^{k}$, respectively. Here, $k$ is the number of GNN layers, usually $k=2$. We leverage the squared quadratic Wasserstein distance (p=2) [vaserstein1969markov] to quantify this similarity, defined as $\mathrm{ED}=W_{2}^{2}(\tilde{H}^{k},H^{k})$.
Pyramid Match Graph Kernel (PMGK) PMGK is a robust measure for graph matching [pmgk], which captures structural similarity across multiple scales. Specifically, we first discretize continuous node embeddings of the recovered and ground-truth graphs into a set of discrete labels via K-means clustering, thereby grouping nodes with similar structural or semantic patterns. We then create a multi-resolution histogram of these labels of each graph and then take a weighted sum of the intersections of these two histograms. In our experiment, PMGK is computed using the grakel library [grakel].
@@ 3.4.4 Performance-level Evaluation
Performance-level evaluation examines whether the recovered unlearned graph is functionally equivalent to the full unlearned graph when processed by the same GNN. This level of evaluation is essential as feature and structural similarities alone cannot guarantee the similar model behavior.
Attack Accuracy (ATT. ACC) Attack accuracy measures whether the recovered graph leads to the same predicted labels as the ground-truth unlearned graph under the same GNN.
$\mathrm{ATT.ACC}=\frac{1}{\left|\mathcal{V}_{u}\right|}\sum_{v\in\mathcal{V}_{u}}\mathbb{I}\Big[\arg\max\tilde{p}_{v}=\arg\max p_{v}\Big]$ (12)
where $\mathbb{I}[\cdot]$ is the indicator function and $p$ is the posterior probability produced by the original GNN $\mathcal{F}_{o}$.
Attack Fidelity (ATT. FID) Attack fidelity quantifies how accurately the recovered graph reproduces the model-specific predictions of the full ground-truth unlearned graph.
$\mathrm{ATT.FID}=\frac{1}{\left|\mathcal{V}_{u}\right|}\sum_{v\in\mathcal{V}_{u}}\mathbb{I}\Big[\arg\max\tilde{p}_{v}=\arg\max\hat{p}_{v}\Big]$ (13)
where $\hat{p}_{v}$ denotes the posterior probability vectors produced by the original GNN on the ground-truth unlearned graph.
[…]
@@ 4 Experimental Validation
In this section, we conduct empirical experiments across a range of GNN variants, graph unlearning methods and graph datasets to answer our three primary research questions. Specifically, we investigate the following key subquestions:
@@
SQ2: How GraphToxin performs for single-node removal?
@@
SQ6: How GraphToxin performs in multiple-node removal?
@@ 4.1.1 Datasets
We conduct experiments on three benchmark graph datasets: Cora, PubMed [cora_pub_data], and Amazon-Photo (Photo) [amazon_photo]. Dataset details are provided in Appendix B.1.
@@ 4.1.2 Unlearning Methods
We consider the following graph unlearning methods. (1) Exact unlearning: We directly retrain the GNNs from scratch (Retraining). (2) Approximate unlearning: We consider two state-of-the-art graph unlearning methods with distinct frameworks: MEGU [MEGU] and ETR [ETR]. MEGU is a learning-based method that employs predictive and unlearning modules to mitigate unlearning effects, while ETR is a training-free method designed to erase the influence of deleted nodes and rectify the collateral damage to their neighbors.
The original training and graph unlearning results across different GNN backbones and graph datasets are detailed in Table 5 in Appendix B.3, which confirm the effectiveness of all considered graph unlearning methods.
@@ 4.1.4 Baselines
To the best of our knowledge, GraphToxin is the first GU-FGRA, there are no established baseline methods for direct comparison. Accordingly, we consider the following alternative baselines: (1) Random Graph (Rand.), which generates random graphs using Gaussian noise without parameter updates. (2) Learning with A Few Epochs (FewE.), which trains GraphToxin for only the first 1% of the total training epochs. This baseline represents a weak attacker with severely limited optimization budget. Direct adaptations of gradient inversion attacks (i.e., gradient matching alone) are evaluated as ablations (Please see Section 4.3.4) rather than standalone baselines, as they represent simplified variants of GraphToxin.
@@ 4.1.5 Implementation
All experiments are implemented on $4$ Nvidia A100100 GPUs. For implementation details, please refer to Appendix B.2.
@@ 4.2 Attack Performance (SQ1 - SQ3)
4.2.1 Necessity of Comprehensive Evaluation (SQ1)
We first examine the necessity of worst-case evaluation. From Table 2, we observe that averaging results under random node removal can obscure meaningful differences in reconstruction difficulty, motivating the need for worst-case evaluation. PubMed exemplifies this problem. Due to its high sparsity and limited label space, a random node removal often produces relatively simple recovery targets, under which even randomly generated graphs may achieve seemingly strong performance-level scores. This effect can artificially inflate average performance and mask the weakness of Rand. on more complex unlearning instances. In contrast, under the worst-case settings, the performance of Rand. significantly decreases, while GraphToxin remains effective, providing a more faithful assessment of reconstruction attacks under challenging unlearning scenarios.
@@ 4.2.1 Necessity of Comprehensive Evaluation (SQ1)
We subsequently motivate a comprehensive evaluation framework for GU-FGRA. Existing attacks commonly rely on RNMSE for assessing feature reconstruction quality [gradient_attack_gnn]. While RNMSE is a standard and necessary metric, it is insufficient when used in isolation. Table 2 shows that even a randomly generated graph may achieve a relatively low RNMSE while performing poorly on performance-level metrics. This discrepancy indicates that RNMSE alone can obscure qualitative differences between the recovered and ground-truth unlearned graphs, which are further amplified by the message-passing paradigm in GNNs. Hence, more sensitive metrics are required to discern subtle variations. Specifically, when RNMSE values are nearly indistinguishable, ED can reveal substantial divergence in the joint feature–structure space. Furthermore, since single-node removal typically induces only subtle parameter changes, reconstruction results across different unlearning methods yield distinct values yet appear nearly identical on the scale of 1e21e\!-\!2. In contrast, PMGK can reflect fine-grained changes that are imperceptible to other metrics.
4.2.2 Effectiveness (SQ2)
As shown in Table 2, GraphToxin consistently and substantially outperforms baselines across all experimental settings on nearly every reconstruction quality and performance-level evaluation metrics. The superiority of GraphToxin generalizes across diverse graph datasets and graph unlearning methods. In addition, GraphToxin’s overall advantage holds under both random and worst-case node removal scenarios, highlighting its robustness. On performance-level metrics, GraphToxin achieves up to $\mathbf{11.14\times}$ improvement over Rand. on Cora, and up to $\mathbf{86.73\times}$ on Photo.